# TorchSISSO: A PyTorch-Based Implementation of the Sure Independence Screening and Sparsifying Operator for Efficient and Interpretable Model Discovery

### v1

@@ 1 Introduction
Additionally, the combinatorial expansion of the feature space may be slow or even infeasible, depending on the available memory. To address this issue, TorchSISSO uses parallel computing and optional GPU acceleration, providing significant computational speed up and scalability of the SISSO method. The remainder of the manuscript is organized as follows: first, we provide a detailed description of the SISSO framework in Section 2, and introduce the proposed toolbox in Section 3. In Section 4, we present performance comparison metrics for the proposed TorchSISSO to the FORTRAN-SISSO implementation. Lastly, we provide concluding remarks in Section 5.
@@ 2 The SISSO Method
The SR problem can be formulated as an empirical risk minimization over a function space $\cal F$. For given target variables $y^{(i)}\in\mathbb{R}$ and feature variables $x^{(i)}\in\mathbb{R}^{d}$ for $i\in\{1,...,N\}$ data points, the SR problem can be defined as [18]
@@
Here, $\mathcal{F}$ consists of all mappings $f:\mathbb{R}^{d}\to\mathbb{R}$ and $f^{\star}$ is the optimal model that produces the lowest average loss $L(\cdot)$, across the training data.
@@ 2.1 Feature Space Expansion
The choice of $\mathcal{F}$ is completely up to the user, however, in general it will contain potentially too many functions to even store in memory. Therefore, SISSO aims to recursively build a set of “expanded features” by applying a set of operators to all possible combinations of features. Let $\boldsymbol{\phi}_{0}=x$ be the initial features and let $\mathcal{O}$ denote the operator set that consists of some number of unary $o[x_{i}]$ and binary $o[x_{i},x_{j}]$ operators. Then, we define the expanded features at level $l\geq 1$ recursively as follows
@@
Note that this is an upper bound since it is possible that some of the combinations are not unique. In the example above, we get $d_{1}\leq 2+\left(\frac{2}{2}\right)(2)(1)=4$, which is exact for the first level since all combinations are unique. For the second level, we get $d_{2}\leq 4+\left(\frac{2}{2}\right)(4)(3)=12$. This bound is larger than the 14 unique combinations shown above because we can exclude, e.g., $x_{1}+x_{2}$ and $x_{1}\times x_{2}$ that would be regenerated when expanding from level 1 to 2.
@@
which grows exponentially with the number of levels $l$ (and the number of binary operators in the operator set). In practice, we can limit this growth by performing dimensional analysis during the expansion process, which restricts certain operators from being applied (e.g., addition and subtraction can only be applied if the features share the same units). However, this does place a strong limit on the maximum expansion level in SISSO – typically needs to be below 4, except in special cases. Also, note that our implementation, described in Section 3, enables the user significant flexibility in their choice of operator set $\mathcal{O}$, which plays a major role on the growth in the feature space.
@@ 4 Numerical Examples
In this section, we compare the performance of TorchSISSO with the original SISSO implementation, referred to as FORTRAN-SISSO, and its derivatives across various test cases, including synthetic equations, challenging scientific benchmarks, and a real-world application in molecular property prediction. All results are based on a single realization of training data generated from the ground-truth equations, potentially corrupted by random observation/measurement noise. However, we found the results to be largely insensitive to the specific data realization. The experiments were run on a computing cluster with two nodes, each equipped with an Intel Xeon Gold 6444Y processor (16 cores) and 512 GB of DDR4 RAM.
@@ 4.3 Interpretable models for molecular property prediction
where $E$ is the redox potential, $E_{\text{anode}}$ is the redox potential of the anode (in this case a Zinc anode), $n$ is the number of moles of electrons transferred, $F$ is Faraday’s constant, and $M_{W}$ is the molecular weight of the molecule. All quantities in (10) are known except for $E$, which can be approximated using density functional theory (DFT). The challenge, however, is that DFT is computationally expensive, making it impractical to scale (10) to millions of candidate molecules. Larger candidate sets are essential when the goal is to discover multiple high-performance molecules. To address this, we construct our training set by sampling data from a literature database presented in [23]. Specifically, we use results for 115 paraquinone molecules as our training set and reserve 1,000 quinone molecules for testing. A crucial step in building molecular property models is featurization, which involves selecting a suitable representation of molecular structure for computational analysis. For this purpose, we use the open-source PaDEL package [24] to compute 1,445 molecular descriptors for each molecule. These descriptors range from basic features, such as atom counts and molecular weight, to more complex graph-based properties. Given the high dimensionality of this problem ($d=1445$), traditional SISSO is not applicable. To manage this, we employ the mutual information (MI) screening approach in TorchSISSO, using the setting initial_screening = ["mi", 0.01] to retain only the top 1% of descriptors by MI value, which reduces the feature set to 11 out of the original 1,445. For comparison, we evaluate TorchSISSO against VS-SISSO [25], an extension of SISSO designed for high-dimensional problems that uses pre-screening. Note that VS-SISSO relies on the original FORTRAN-SISSO code for backend computations, making it a useful benchmark for our case study.
@@
The training and testing results for both TorchSISSO and VS-SISSO are shown in Figure 2. We see that both approaches are able to obtain good training performance, with TorchSISSO and VS-SISSO achieving $R^{2}$ values of 0.985 and 0.936, respectively. However, we see a bigger difference on the test data wherein TorchSISSO and VS-SISSO achieve $R^{2}$ values of 0.932 and 0.604, respectively. In particular, VS-SISSO shows a significant drop in performance for specific energy values below 0.75 where it clearly has a biased over-prediction in this range. TorchSISSO, on the other hand, has a much tighter parity plot throughout the full range of specific energy values, implying it has learned an equation that generalizes much better beyond than the training dataset.
@@
Specific $\text{Specific Energy}\approx 144.14676\left(\frac{P_{GH}+\lambda_{M}}{M_{W}\times P_{GH}}\right)+0.06388,$ (11)
@@
where $P_{GH}$ is the solute gas-hexadecane partition coefficient, $\lambda_{M}$ is the largest absolute eigenvalue of the Burden modified matrix weighted by relative mass, and $M_{W}$ is molecular weight. One interesting thing to notice right away is that (11) has exactly the same $M_{W}$ term in the denominator as (10) – we emphasize that this structure was not imposed during the training process, but was uncovered directly from the data. Although the other two features $P_{GH}$ and $\lambda_{M}$ are not quite as intuitive, they do carry physical significance. For example, $P_{GH}$ provides a measure of how a molecule interacts with solvents, which can impact the electronic properties (such as redox potential). Despite starting with a large and complex set of potential descriptors, TorchSISSO was able to pinpoint a compact, interpretable equation that relies on just three fundamental molecular features, combined in a simple form, to achieve high predictive accuracy on both the training and test sets. Furthermore, the specific implementation choices clearly result in an improvement over the state-of-the-art VS-SISSO code for at least this real-world example.

### v2

@@ 1 Introduction
Additionally, the combinatorial expansion of the feature space may be slow or even infeasible, depending on the available memory. To address this issue, TorchSISSO uses parallel computing and optional GPU acceleration, providing significant computational speed up and scalability of the SISSO method. The remainder of the manuscript is organized as follows: first, we provide a detailed description of the SISSO framework in Section 2, and introduce the proposed toolbox in Section 3. In Section 4, we present performance comparison metrics for the proposed TorchSISSO to the FORTRAN-SISSO package. Lastly, we provide concluding remarks in Section 5.
@@ 2 The SISSO Method
The SR problem can be formulated as an empirical risk minimization over a function space $\cal F$. For given target variables $y^{(i)}\in\mathbb{R}$ and (base) feature variables $x^{(i)}\in\mathbb{R}^{d}$ for $i\in\{1,...,N\}$ data points, the SR problem can be defined as [18]
@@
The target (also known as outcome or response) represents the dependent variable that is assumed to be a deterministic transformation of a set of base features (also known as inputs) that are the independent variables in the problem. Here, $\mathcal{F}$ consists of all possible mappings $f:\mathbb{R}^{d}\to\mathbb{R}$ (needs to be specified by the user) and $f^{\star}$ is the optimal model that produces the lowest average loss $L(\cdot)$, across the training data.
@@ 2.1 Feature Space Expansion
The choice of $\mathcal{F}$ is completely up to the user, however, in general it will contain potentially too many functions to even store in memory. Therefore, SISSO aims to recursively build a set of “expanded features” by applying a set of operators to all possible combinations of features. We highlight that different choices of operators and base features (inputs) will lead to different constructions of $\mathcal{F}$. The goal of SISSO is then to efficiently build and search over a large number of expanded features for modeling the target variable. Let $\boldsymbol{\phi}_{0}=x$ be the initial features and let $\mathcal{O}$ denote the operator set that consists of some number of unary $o[x_{i}]$ and binary $o[x_{i},x_{j}]$ operators. Then, we define the expanded features at level $l\geq 1$ recursively as follows
@@
Note that this is an upper bound since it is possible that some of the combinations are not unique. In the example above, we get $d_{1}\leq 2+\left(\frac{2}{2}\right)(2)(1)=4$, which is exact for the first level since all combinations are unique. For the second level, we get $d_{2}\leq 4+\left(\frac{2}{2}\right)(4)(3)=16$. This bound is larger than the 14 unique combinations shown above because we can exclude, e.g., $x_{1}+x_{2}$ and $x_{1}\times x_{2}$ that would be regenerated when expanding from level 1 to 2.
@@
which grows exponentially with the number of levels $l$ (and the number of binary operators in the operator set). In practice, we can limit this growth by performing dimensional analysis during the expansion process, which restricts certain operators from being applied (e.g., addition and subtraction can only be applied if the features share the same units). However, this does place a strong limit on the maximum expansion level in SISSO – typically needs to be below 4, except in special cases. Also, note that our implementation, described in Section 3, enables the user significant flexibility in their choice of operator set $\mathcal{O}$, which plays a major role on the growth in the feature space. In practice, one may treat $l$ as a tunable hyperparameter, as the best choice will depend on the problem. Given the steep increase in cost as $l$ increases, we recommend a greedy approach: start with small $l$ (likely $l=1$) and incrementally increase it until model accuracy meets the requirements of the application. If one is unable to find a sufficiently good model with $l\leq 3$, it is recommended to revisit the choice of operators included in $\mathcal{O}$ before attempting even higher expansion levels.
@@ 2.3 Sparsifying Operator
2.4 Impact of Noise and Data Distribution
The SISSO framework assumes that the target variable is a (sparse) linear combination of the expanded features, with observations potentially corrupted by additive random noise:
$y^{(i)}=\boldsymbol{\phi}(x^{(i)})^{\top}\boldsymbol{c}+\varepsilon^{(i)},\quad i=1,\ldots,N,$ (9)
where each $\varepsilon^{(i)}$ is an unobserved error term assumed to follow a given probability distribution. The assumed distribution of $\varepsilon^{(i)}$ influences the choice of loss function in (1). Currently, our implementation focuses on the commonly used “least squares” formulation, derived from maximum likelihood estimation (MLE) under the assumption that the noise $\{\varepsilon^{(i)}\}_{i=1}^{N}\sim\mathcal{N}(0,\sigma^{2})$ follows an i.i.d. zero-mean Gaussian distribution.
In cases where the noise vector $\boldsymbol{\varepsilon}=(\varepsilon^{(1)},\ldots,\varepsilon^{(N)})$ instead follows a more general multivariate Gaussian distribution, $\boldsymbol{\varepsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$, a whitening transformation can be applied to map the problem to a space where the noise terms are i.i.d., provided $\boldsymbol{\Sigma}$ is known. Specifically, we compute a weight matrix $\boldsymbol{W}=\boldsymbol{\Sigma}^{-1/2}$ and define transformed targets and features, $\tilde{\boldsymbol{y}}=\boldsymbol{W}\boldsymbol{y}$ and $\tilde{\boldsymbol{\Phi}}=\boldsymbol{W}\boldsymbol{\Phi}$, satisfying the i.i.d. noise assumption. After SISSO identifies a model in the transformed space, this result can be converted back to the original scale as needed for prediction and interpretation.
The level of noise in the observations can significantly affect the ability to accurately learn the governing equations. In practice, given finite data, the best-fit models will depend on interrelated factors such as the number of data points, the noise level, the structure of the ground-truth equations, and the range of the observed data. Since these dependencies are often difficult to understand a priori, standard practice in machine learning emphasizes extensive validation and testing. Validation on held-out test data is one straightforward approach, but for low-data settings, $k$-fold cross-validation is generally preferred as it leverages the full dataset while providing an estimate of performance variation across subsets of the data. Regardless of the validation method, users should remain cautious of the “best” model identified by SISSO or any other algorithm until it has been thoroughly evaluated in tasks that reflect its intended application. Robustness to noise can also be enhanced by ensembling techniques, such as bootstrap aggregation (bagging) [19], which has been shown to improve model resilience, as demonstrated with the SINDy algorithm in [20]. Similar approaches may be applied to SISSO, a direction we intend to pursue in future work.
Additionally, the distribution of the input variables (base features $x$) plays a critical role in the model learned. For instance, if all measurements are concentrated within a narrow range in the input space, the data will contain limited information about the overall target variable distribution, and consequently, learning a broadly accurate model becomes nearly impossible. Section 4.4 provides an illustrative example showing how a model can achieve high accuracy within a localized region of the input space, highlighting the importance of validation, particularly when extrapolating beyond the observed data range.
@@ 4 Numerical Examples
In this section, we compare the performance of TorchSISSO with the original SISSO implementation, referred to as FORTRAN-SISSO, and its derivatives across various test cases, including synthetic equations, challenging scientific benchmarks, and a real-world application in molecular property prediction. All results are based on a single realization of training data generated from the ground-truth equations, potentially corrupted by random observation/measurement noise. However, we found the results to be largely insensitive to the specific data realization. The experiments were run on a computing cluster with two nodes, each equipped with an Intel Xeon Gold 6444Y processor (16 cores) and 512 GB of DDR4 RAM. All the experiments, except those in Section 4.5, are run on a CPU to provide a fair comparison with FORTRAN-SISSO, which currently does not provide GPU support.
@@ 4.3 Interpretable Models for Molecular Property Prediction
where $E$ is the redox potential, $E_{\text{anode}}$ is the redox potential of the anode (in this case a Zinc anode), $n$ is the number of moles of electrons transferred, $F$ is Faraday’s constant, and $M_{W}$ is the molecular weight of the molecule. All quantities in (11) are known except for $E$, which can be approximated using density functional theory (DFT). The challenge, however, is that DFT is computationally expensive, making it impractical to scale (11) to millions of candidate molecules. Larger candidate sets are essential when the goal is to discover multiple high-performance molecules. To address this, we construct our training set by sampling data from a literature database presented in [25]. Specifically, we use results for 115 paraquinone molecules as our training set and reserve 1,000 quinone molecules for testing. A crucial step in building molecular property models is featurization, which involves selecting a suitable representation of molecular structure for computational analysis. For this purpose, we use the open-source PaDEL package [26, 27] to compute 1,875 molecular descriptors for each molecule. These descriptors range from basic features, such as atom counts and molecular weight, to more complex graph-based properties. There are a total of 1444 one- and two-dimensional descriptors and 431 three-dimensional descriptors. Before moving to the modeling phase, we first compute the variance of all descriptors over the training set and remove those below a small threshold, leaving us with 1445 possible descriptors. Given the high dimensionality of this problem ($d=1445$), traditional SISSO is not applicable. To manage this, we employ the mutual information (MI) screening approach in TorchSISSO, using the setting initial_screening = ["mi", 0.01] to retain only the top 1% of descriptors by MI value, which reduces the feature set to 11 out of the original 1,445. For comparison, we evaluate TorchSISSO against VS-SISSO [28], an extension of SISSO designed for high-dimensional problems that uses pre-screening. Note that VS-SISSO relies on the original FORTRAN-SISSO code for backend computations, making it a useful benchmark for our case study. No additional features were pre-screened beyond those with near zero variance, so VS-SISSO has access to the original 1,445 descriptors. We used the default settings for VS-SISSO, with two minor modifications: (i) we set the maximum number of base features n_max to be 11 to match that allowed for the MI method and to ensure the computational time was reasonable and (ii) we increased the maximum number of iterations nstep_max to 200 to allow increased exploration of the feature space.
@@
The training and testing results for both TorchSISSO and VS-SISSO are shown in Figure 2. We see that both approaches are able to obtain good training performance, with TorchSISSO and VS-SISSO achieving $R^{2}$ values of 0.985 and 0.936, respectively. However, we see a bigger difference on the test data wherein TorchSISSO and VS-SISSO achieve $R^{2}$ values of 0.932 and 0.604, respectively. In particular, VS-SISSO shows a significant drop in performance for specific energy values below 0.75 where it clearly has a biased over-prediction in this range. TorchSISSO, on the other hand, has a much tighter parity plot throughout the full range of specific energy values, implying it has learned an equation that generalizes much better beyond than the training dataset. The CPU time required by TorchSISSO was just 14.3 seconds, representing an over 16-fold reduction compared to the 237.1 seconds needed for VS-SISSO, while also yielding a model with better predictive accuracy.
@@
Specific $\text{Specific Energy}\approx 144.14676\left(\frac{P_{GH}+\lambda_{M4}}{M_{W}\times P_{GH}}\right)+0.06388,$ (12)
@@
where $P_{GH}$ is the solute gas-hexadecane partition coefficient, $\lambda_{M4}$ is the largest absolute eigenvalue of the Burden modified matrix weighted by relative mass with modification parameter $n=4$, and $M_{W}$ is molecular weight. One interesting thing to notice right away is that (12) has exactly the same $M_{W}$ term in the denominator as (11) – we emphasize that this structure was not imposed during the training process, but was uncovered directly from the data. Although the other two features $P_{GH}$ and $\lambda_{M4}$ are not quite as intuitive, they do carry physical significance. For example, $P_{GH}$ provides a measure of how a molecule interacts with solvents, which can impact the electronic properties (such as redox potential). Despite starting with a large and complex set of potential descriptors, TorchSISSO was able to pinpoint a compact, interpretable equation that relies on just three fundamental molecular features, combined in a simple form, to achieve high predictive accuracy on both the training and test sets. Furthermore, the specific implementation choices clearly result in an improvement over the state-of-the-art VS-SISSO code for at least this real-world example.
Note that the best (lowest error) equation found by VS-SISSO can be expressed as:
Specific $\text{Specific Energy}\approx 1.03252\left(\frac{\text{ATS3i}}{\lambda_{M1}}\right)+0.5533907820,$ (13)
where ATS3i is the Broto-Moreau autocorrelation with lag 3 weighted by the first ionization potential and $\lambda_{M1}$ is the largest absolute eigenvalue of the Burden modified matrix weighted by relative mass with modification parameter $n=1$. Interestingly, VS-SISSO found a different Burden eigenvalue descriptor than TorchSISSO, which appears in the denominator as opposed to the numerator in the overall learned descriptor. VS-SISSO also does not identify the importance of molecular weight in the expression, which might be the source of the biased predictions for molecules with low specific energy values.
4.4 Illustration of Potential Challenges with Extrapolation
In this section, we examine how the distribution of input data affects SISSO’s performance, particularly when extrapolating beyond a narrow data range. We aim to highlight challenges that arise from limited training distributions, an issue driven by data quality rather than the SISSO method itself. As an example, we use an unknown ground-truth equation based on a modified Arrhenius form:
$k=2.37\sqrt{T}e^{-\frac{E_{a}}{RT}},$ (14)
where $k$ is the rate constant, $T$ the temperature, $E_{a}$ the activation energy, and $R$ the ideal gas constant. We simulate noisy observations from (14), where $y=k+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,0.1)$, across various temperature ranges. For simplicity, we set $E_{a}=185$ kJ/mol and $R=8.314\times 10^{-3}$ kJ/mol-K, aiming to recover (14) using SISSO. We assume that $E_{a}$ is known such that our input features include $E_{a}$ and $RT$, and the operator set $\mathcal{O}=\{\sqrt{(\cdot)},(\cdot)+(\cdot),\exp(-(\cdot)),(\cdot)/(\cdot)\}$.
First, we train SISSO on 100 randomly generated samples from a “training range” of $T\in[800,900]$ K. The resulting SISSO model after training is:
$k\approx 1385.95e^{-\frac{E_{a}}{RT}}-1282.82,$ (15)
which achieves a training $R^{2}$ of $\sim 0.99$. We validate this model on two extrapolation datasets: one in a range $T\in[710,795]$ K near that of the training data and another in a farther range of $T\in[600,700]$ K. Figure 3 (left) shows the parity plots, indicating that despite high accuracy within the training range, the model deviates significantly from ground truth as $T$ moves outside this range. This is expected, as the learned model structure does not fully capture the temperature dependency of the frequency factor. The close alignment of (15) to observed data within the training range masks its limitations, illustrating the need for rigorous testing across varied inputs, especially outside the training range.
Upon observing these deviations, one practical approach is to augment the training set with data that spans a wider temperature range. Doing so here enables SISSO to recover exactly the ground-truth structure in (14). Figure 3 (right) shows the parity plot for a model trained on an 80%/20% training/validation split over this expanded range. The resulting SISSO model is:
$k\approx 2.3723\sqrt{T}e^{-\frac{E_{a}}{RT}}-0.0752,$ (16)
which correctly captures the frequency factor’s temperature dependence due to the improved data distribution.
In summary, while it may not always be possible to recover the exact structure of the ground-truth equation, this limitation does not prevent the effective use of learned models. Through systematic validation, we can rigorously assess model performance and identify any limitations in predictive capability. Additionally, by intentionally holding out data points near the extremes of the input space, we can further probe the model’s extrapolation capacity. Such testing methods enable us to pinpoint areas where the model may struggle. When models demonstrate consistent performance across these held-out extremes, we can be more confident in their ability to generalize beyond the training range. This approach allows us to leverage these models with greater assurance, even when the true underlying structure remains only partially known.
Figure 3: (Left) Parity plot showing training and validation results for a SISSO model trained on data collected over a limited temperature range, with validation data spanning temperatures both near and far outside this range. (Right) Parity plot showing training and validation results for a SISSO model trained on data spanning the full temperature range, with an 80/20 train/validation split. This comparison illustrates the impact of training data distribution on model generalizability across a broader temperature spectrum.
4.5 Computational Time Comparison on CPU and GPU Hardware
To further demonstrate the advantages of TorchSISSO on hardware accelerators, specifically GPUs, we compare runtime performance between TorchSISSO and FORTRAN-SISSO across several computing configurations. While the default version of FORTRAN-SISSO is limited to single-core processors, a recent update introduces multi-core support. However, due to the complex installation requirements for multi-core usage, we could not implement this on our cluster, so this configuration is excluded from the comparisons. A key advantage of TorchSISSO is its versatility: it can be readily installed and executed on single-core and multi-core CPUs, as well as on GPUs, and we examine all these configurations in this section. We used an NVIDIA A100 Tensor Core GPU with 40 GB of RAM on our computing cluster, capping CPU memory to 40 GB to ensure a consistent maximum RAM across all tests.
To assess runtime performance on more complex regression tasks, we constructed three four-term models: (i) $x_{1}^{4}+x_{2}^{3}+x_{3}^{2}+x_{4}$; (ii) $x_{1}^{5}+x_{2}^{4}+x_{3}^{2}+x_{4}$; and (iii) $x_{1}^{3}+x_{2}^{2}+x_{3}+\sin(x_{4})$. For each model, 100 data points were generated, with each input component sampled uniformly over $[1,5]$. Execution times for FORTRAN-SISSO (single CPU core) and TorchSISSO (single- and multi-core CPU, and GPU) as a function of the parameter $k$ are illustrated in Figure 4. Error bars reflect minimum and maximum times across the three test cases. As expected, training times increase with $k$, which determines the number of features selected per term for regression and, hence, the number of models fit. Across all cases, GPU runtime is the shortest, with FORTRAN-SISSO as the second fastest, likely due to advantages inherent in optimized Fortran compilers for scientific computing. It is worth noting that TorchSISSO was slightly faster than FORTRAN-SISSO on most of the previous benchmarks (Section 4.14.2) since the equations of interest were not as complicated as that considered in this section and we did not need to cap the CPU memory to fairly compare against the GPU. Moreover, we used default, non-optimized settings for TorchSISSO in all of our runs, so further optimization has the potential to yield even greater speed improvements. Despite the lack of any systematic optimization, the GPU results show a substantial (2-4x) reduction in time compared to the optimized Fortran implementation.
Figure 4: Computational time versus parameter $k$ for running TorchSISSO and FORTRAN-SISSO on different hardware. Note that $k$ controls how many models must be fit according to (6)–(8).
Finally, we assess performance for ensemble model training. Using the first four-term model, we trained 30 different models by partitioning the data into 30 subsets of 100 samples each (fixing $k=20$ to its default value). In the torch framework, all training tasks can be dispatched simultaneously to the GPU, enabling substantial time savings compared to sequential CPU and Fortran implementations, as shown in Table 3.
Table 3: Computational times (seconds) for training an ensemble of 30 models each with 100 datapoints for a four term ground-truth expression using FORTRAN-SISSO and TorchSISSO under different hardware configurations.
FORTRAN-SISSO TorchSISSO (single-CPU) TorchSISSO (multi-CPU) TorchSISSO (GPU)
182.43 321.87 287.34 51.32